OpenStudy (anonymous):

The linear approximation to (1/(2-x)^(1/2) at x=0 is y=???? Help out with this problem??

OpenStudy (anonymous):

\[f(x)=\frac{1}{\sqrt{2-x}},f(0)=\frac{1}{\sqrt{2}}\] \[f'(x)=\frac{1}{2\sqrt{(2-x)^3}},f'(0)=\frac{1}{2\sqrt{8}}\]\]

OpenStudy (anonymous):

equation for the line is \[y-\frac{1}{\sqrt{2}}=\frac{1}{4\sqrt{2}}x\] or \[y=\frac{1}{4\sqrt{2}}x+\frac{1}{\sqrt{2}}\] feel free to rationalize the denominators